# DesignRepair: Dual-Stream Design Guideline-Aware Frontend Repair with Large Language Models

### v1

@@ I Introduction
The advent of advanced large language models (LLMs), such as GPT-4 [1] and Meta’s Llama [2], has catalyzed transformative changes across various fields like legal [3, 4, 5, 6], finance [7], and software development [8], unlocking a plethora of new possibilities. For frontend software development, commercial and open-source tools, such as Vercel’s V0[9], Imagica[10] and OpenV0 [11], have attracted considerable interest using the capabilities of LLMs. These applications adeptly transform user-provided textual or visual prompts into concrete, well-structured high-fidelity user interfaces, along with their associated front-end code, providing substantial support to designers and developers. To augment the quality of the generated UI and its accompanying code, these tools commonly integrate with established UI design libraries (e.g., HTML, CSS, JS frameworks, and libraries), including popular ones like React [12], Shadcn [13] and Tailwind CSS [14]. These libraries serve as foundational knowledge bases, empowering these tools to generate aesthetically pleasing and technically sound designs.
Despite their advanced capabilities, these tools continue to face two significant challenges. Firstly, while they are adept at converting prompts into well-structured UI elements, their compliance with established design principles, such as material design [15], often falls short (see Section IV). This inadequacy is critical, as stringent adherence to these principles is fundamental to crafting effective, accessible, and user-friendly interfaces. Such high-quality interfaces play a pivotal role in setting an application apart from its competitors, boosting user downloads, reducing user complaints, and enhancing user retention [16, 17, 18, 19, 20, 21]. Secondly, these tools exhibit limitations in the repair and refinement of generated designs. While they allow for additional prompts for UI adjustments, they often rely on users to manually identify and describe needed changes. This process can be cumbersome, requiring extensive labeling and in-depth knowledge [18, 19, 22]. Fig. 1(a) shows an example. When prompted with “URL shorten Chrome extension,” Vercel’s V0 successfully generated an appealing user interface. Nonetheless, upon closer examination, it exhibited five distinct design issues related to contrast and color in text and button elements, marked in red.
@@
Many studies have been proposed to assess different facets of UI design development to ensure its quality, encompassing accessibility [23, 24, 25, 26, 27, 28], tapability [29], animation effects [18], and broader design principles and concerns [19, 30, 16]. Industrial tools like Playwright [23] and Google Lighthouse [24] primarily address accessibility concerns, such as text contrast for readability. Swearing et al.[29] utilized a classifier-based technique to identify discrepancies between the perceived and intended tappability of UI elements. Yang et al. [19] introduced UIs-Hunter, a rule-based method for detecting deviations from material design principles in UIs. Despite these advancements, they either tend to focus on specific, isolated issues within UI design, or they encounter challenges in generalizing to a broader spectrum of design aspects. Additionally, while some solutions [31, 21] have been proposed to rectify detected issues, they too tend to concentrate on limited aspects of design principles. The narrow scope of rule-based methods reveals that encoding every guideline for different platforms is labor-intensive and only effective in specific environments.
@@
We conducted an evaluation to measure the effectiveness and usefulness of DesignRepair. To carry this, we collected 196 design issues from AI-generated frontend code and 115 from GitHub projects consisting of UI design and its frontend code. Our evaluation shows that DesignRepair is able to identify and repair design issues, with a recall of 89.3% and precision of 86.6% for AI-generated issues and a recall of 85.2% and precision of 90.7% for GitHub issues. Additionally, our experiments on each step highlights the effectiveness of the strategies we implemented. Finally, a user study with 26 participants further confirms the high quality of our repairs.
@@
  • We systematically analyzed Material Design guidelines and systematically converted them into structured knowledge bases, i.e., low level component knowledge base and high-level system design knowledge base, enabling efficient and effective utilization of knowledge in design quality assurance processes.
  • We introduced DesignRepair, a novel dual-stream knowledge-driven, LLM-based method, uniquely designed to systematically detect and repair design quality issues in frontend code. This approach simultaneously considers both the source code and the user-perceived rendered page, ensuring outstanding performance in design quality enhancement.
@@ II-A Material Design Guidelines
Material Design 3 [15] is a broader design system built and supported by Google designers and developers, and can be implemented across Android, iOS, and web platforms. Its latest version, Material Design 3, provides guidelines and examples to support important practices of user interface, ensuring personal, adaptive, and expressive experiences. It covers topics ranging from fine-grained design guidelines such as the application of individual components and color usage for specific meanings, to the high-level composition of components and common design patterns.
@@ III Approach
For the offline knowledge base construction phase (Fig. 3-A), we built a two-part knowledge base (KB): a low-level Component Knowledge Base (KB-Comp) and a high-level System Design Knowledge Base (KB-System). This knowledge base functions as a domain expert, offering guidance for addressing potential UI issues. Given the frontend code and rendered page, we enter the second phase, where we use a parallel dual-pipe method to extract the used components and their corresponding property groups (as shown in Fig. 3-B). Finally, we implement a knowledge-driven, LLM-based repair method (Fig. 3-C). This approach allows us to meticulously analyze and repair issues concurrently. By employing a divide and conquer strategy, we tackle each component/property group individually before synthesizing the repairs. This ensures a cohesive, optimized final output, achieved through a thorough and scrutinized repair process.
@@ III-A Knowledge Base Construction
As introduced in Section II-A, we utilise Material Design 3 as our foundational knowledge base. By systematically analyzing the guidelines for efficient knowledge retrieval, we restructured them into two parts: a low-level Component Knowledge Base (KB-Comp) and a high-level System Design Knowledge Base (KB-System). KB-Comp provides detailed guidance for atomic-level elements, while KB-System addresses the broader aspects of aesthetic and functional harmony in interface design. For example, a guideline from KB-Comp might ensure a button is sized appropriately for accessibility purpose, while a guideline from KB-System would focus on how this button fits into the overall page layout and flow. Within these knowledge bases, we further classified the information into soft constraints (i.e., recommended guidelines) and hard constraints (i.e., do/don’t guidelines), catering to different levels of design concerns. For instance, a hard constraint may specify the minimum text size for readability, whereas a soft constraint could suggest font styles that align with the brand’s aesthetic.
@@ III-A1 Component knowledge base
We categorised the component knowledge base (KB-Comp) at the component level, drawing from Components section in the Material Design (see Section II-A). Each component is linked with a range of guidelines. In total, we obtained 24 component types. Fig. 4(a) illustrates the number of guidelines for each component type and the distribution of various UI/UX design aspects (like usage, responsive layout) that these guidelines address. The distribution of guidelines across these component types demonstrates a notable imbalance, showing a long-tail distribution pattern. For instance, buttons and navigation components are the most abundant, with 109 and 95 entities respectively, while checkboxes are the least represented, with only 6 entities. To facilitate a better understanding, Fig. 5 (a) shows two hard constraints and two soft constraints regarding the button component. In this example, the first hard constraint aligns with the guidelines shown in Fig. 2(b). Similarly, the first soft constraint correlates with the guidelines in Fig. 2(a).
@@
In conclusion, based on our detailed knowledge base, we propose Knowledge-driven LLM method, which can effectively manage the long-tail distribution of design guidelines, ensuring an equitable application of these principles across all component types.
@@ III-A2 System Design knowledge base
To this end, we reformulated high-level design guidelines, i.e., Foundations and Styles sections (see Section II-A, into a structured system design knowledge base (KB-System). In total, We identified 12 system design aspects (e.g., typography) from Foundations and Styles, summarised 7 properties (e.g., clickable and spacing) and 15 mapping relations. For example, ‘role’ and ‘utilities’ from Styles Color and accessibility ‘contrast’ from Foundations Color are both connected to the property ‘color’. These relations are crucial for marking guidelines, aiding LLMs in understanding the intricate relationships between design elements and their practical applications in a web development context. Based on this, we reconstructed the design guidelines into triples of $<$system design aspect, property, relation$>$. Fig. 5 (b) shows two examples of the system design knowledge base.
Foundations, shown in red, have 7 key aspects including Flow, Layout, Implementation, Structure, Label, Text, and Color. For instance, the ‘Flow’ aspect relates to how users navigate through an interface, while ‘Color’ pertains to both aesthetic appeal and accessibility. In total, we identified 118 knowledge entities for Foundations.
Styles, depicted in blue, consist of 5 key aspects: Typography, Elevation, Icons, Color, and Shape. These elements craft the UI’s visual identity, adaptable through Material themes. In our work, we structured 80 knowledge entities related to Styles, with Icons specifically linked to icon buttons and Shapes applicable across all component types.
Property Groups (golden color) include 7 categories: Group, Clickable, Spacing, Platform, Label, Text, and Color. These are actionable property levels that we map to high-level guidelines for effective detection and repair.
@@ III-B1 Components Extraction
The first aspect of our strategy is to identify a list of component types within the Page Under Test (PUT) using GPT-4. This initial step is pivotal for the efficient retrieval of relevant component knowledge from KB-Comp in subsequent repair stages.
@@ III-C1 Relevant Knowledge Retrieval
(a) Component Mapping and Knowledge Retrieval: To map extracted web page components to the component names in our KB-Comp for alignment, we use LLMs for reasoning with the prompt $P_{\text{map\_kb}}$. This approach allows us to handle the diversity in terminology used across different front-end libraries such as Shadcn[13], Tailwind CSS[14] without defining any exhaustive rule-based mapping function. For instance, a ‘navigation bar’ component might be referred to as ‘Navbars’ in Tailwind CSS, while in Shadcn UI it could be labeled as ‘Navigation Menu’. Such variability in naming conventions can be effectively addressed by LLMs without the need to define an exhaustive mapping function. The mapping process is represented by:
@@ III-C3 Overall Repair
The Overall Repair phase is the culmination of our knowledge-driven repair process. In this final step, we integrate the repair suggestions obtained from both the component-level and property-group analyses. This integration ensures a holistic repair of the frontend code.
Formally, given the original PUT code $In_{\text{PUT}}$ and the set of partially repaired code snippets $C3$, our objective is to consolidate these inputs to produce the final, fully repaired page. To achieve this, we employ prompt with $P_{\text{all}}$, which is designed to orchestrate the overall repair process:
@@
The $P_{\text{all}}$ prompt effectively merges the various repair suggestions, taking into account any potential overlaps or conflicts between the proposed fixes. The goal is to ensure that the final output, represented as Repaired_page, is not only free from identified design flaws but also aligns with the established design guidelines and usability standards. Due to the length limit of the article, all prompts can be obtained on GitHub  [^1] https://github.com/DesignRepair2024/DesignRepair2024.
@@ IV Experiments
  • RQ1: How effective is our approach in detecting violations according to Design Guidelines compare to baseline?
  • RQ2: What is the contribution of each key repair strategy to the overall performance of our approach?
  • RQ3: What is the perceived quality of the fixes generated by our approach from a user perspective?
@@ IV-A Subjects
We conducted two experimental studies to evaluate the performance of our frontend code repair approach on both generated code from Vercel’s V0 and real projects from GitHub.
@@ IV-A1 Vercel’s V0 Projects
Our research focuses on examining code and the rendered pages generated by LLM tools, specifically targeting adherence to design guidelines and enabling code-level adjustments. Our work aims to assist developers with limited knowledge of UI design guidelines who utilize LLM-generated UI mock-ups as starting points. Vercel’s V0 [9] generated projects were chosen for this purpose, as it is a leading LLM tool that attracted 100K users within three weeks of its release [32]. Additionally, Vercel’s V0 is widely used to assist in frontend development, with 175K results on GitHub as of August 2023.
@@ IV-A2 Github Projects
To further evaluate the effectiveness of our method in real-world projects, we searched for “top frontend projects” and “frontend example projects” on GitHub and Google, identifying the top 6 popular projects. These projects represent various real-world website uses, ranging from commercial product frontend websites to blogs, with star counts ranging from 1.1k to 5.6k. Four of these projects are still active, while two have been archived. Their popularity and representativeness have proven their usefulness.
We randomly selected 66 files out of 577 files from these projects, ensuring diverse functionality. These projects utilize a variety of representative packages, including React [12], Next.js [33], Material-UI [34], and TypeScript [35]. Our LLM-based method provides the capability for detection and repair across these different technologies.
Following the same protocol for recognizing guideline violations as we did with Vercel’s V0, we identified 1,793 components/property groups across these projects, of which 115 were violated.
@@ IV-A3 Subject Analysis
For Vercel’s V0 projects, there were 33.45 per file components/property group and 9.8 guideline violations per project. Of these, 149 (76%) were component-related and 47 (24%) were property-related. Additionally, 27 (14%) were soft constraint violations, while 169 (86%) were hard constraint violations. For GitHub projects, there were 27 per file components/property group and 1.74 guideline violations per project. Among these, 60 (52%) were component-related and 55 (48%) were property-related. Soft constraint violations accounted for 15 (13%), and hard constraint violations accounted for 100 (87%).
Firstly, the generated frontend code from Vercel’s V0 was found to have quality comparable to real GitHub projects (33.45 vs. 27 components/property group per file). This demonstrates the practicality of using LLM-generated code in real frontend development. Secondly, Vercel’s V0 projects had a higher percentage of component-related issues (52% more) compared to property-level issues. Real-world projects, although generally better designed, still had minor issues such as inconsistent padding or missing alt-text. Thirdly, the higher number of hard constraint violations is due to the difficulty in identifying soft constraint violations and the overall high quality of both Vercel’s V0 and GitHub projects. Both datasets exhibited a few obvious design flaws such as poor color usage or layout issues. Comparatively, GitHub projects had fewer violations per file than Vercel’s V0 projects (1.74 vs. 9.8), indicating that LLM-generated results require more automated quality assurance to ensure code quality. In conclusion, human-generated issues differ significantly from AI-generated ones, making it valuable to analyze them separately for better insights.
IV-B RQ1: How effective is our approach in detecting violations according to Design Guidelines compare to baseline?

### v2

@@ I Introduction
The emergence of advanced large language models (LLMs), such as GPT-4 [1] and Meta’s Llama [2], has catalyzed transformative changes across various fields, including legal [3, 4, 5, 6], finance [7], and software development [8], unlocking tremendous new possibilities. For frontend software development, commercial and open-source tools, such as Vercel’s V0[9], Imagica[10] and OpenV0 [11], have attracted considerable interest using the capabilities of LLMs. These applications adeptly transform user-provided textual or visual prompts into concrete, well-structured high-fidelity user interfaces, along with their associated front-end code, providing substantial support to designers and developers. To augment the quality of the generated UI and its accompanying code, these tools commonly integrate with established UI design libraries (e.g., HTML, CSS, JS frameworks, and libraries), including popular ones like React [12], Shadcn [13] and Tailwind CSS [14]. These libraries serve as foundational knowledge bases, empowering these tools to generate aesthetically pleasing and technically sound designs.
Despite their advanced capabilities, these tools continue to face two significant challenges. Firstly, while they are adept at converting prompts into well-structured UI elements, their compliance with established design principles, such as material design [15], often falls short (see Section IV). This inadequacy is critical, as strict adherence to these principles is fundamental to crafting effective, accessible, and user-friendly interfaces. Such high-quality interfaces play a vital role in setting an application apart from its competitors, boosting user downloads, reducing user complaints, and enhancing user retention [16, 17, 18, 19, 20, 21]. Secondly, these tools exhibit limitations in the repair and refinement of generated designs. While they allow for additional prompts for UI adjustments, they often rely on users to manually identify and describe needed changes. This process can be cumbersome, requiring extensive labeling and in-depth knowledge [18, 19, 22]. Fig. 1(a) shows an example. When prompted with “URL shorten Chrome extension,” Vercel’s V0 successfully generated an appealing user interface. Nonetheless, upon closer examination, it exhibited five distinct design issues related to contrast and color in text and button elements, marked in red.
@@
Many studies have been proposed to assess different facets of UI design development to ensure its quality, encompassing accessibility [23, 24, 25, 26, 27, 28], tapability [29], animation effects [18], and broader design principles and concerns [19, 30, 16]. Industrial tools like Playwright [23] and Google Lighthouse [24] primarily address accessibility concerns, such as text contrast for readability. Swearing et al.[29] utilized a classifier-based technique to identify discrepancies between the perceived and intended tappability of UI elements. Yang et al. [19] introduced UIs-Hunter, a rule-based method for detecting deviations from material design principles in UIs. Despite these advancements, they either tend to focus on specific, isolated issues within UI design, or they encounter challenges in generalizing to a broader spectrum of design aspects. Moreover, rule-based methods require extensive manual efforts which limits their usability and generalisability.
@@
We conducted an evaluation to measure the effectiveness and usefulness of DesignRepair. We collected 196 design issues from AI-generated frontend code and 115 from GitHub projects consisting of UIs and their frontend code. Our evaluation shows that DesignRepair effectively identifies and repairs design issues, achieving a recall of 89.3% and precision of 86.6% for AI-generated issues, and a recall of 85.2% with precision of 90.7% for GitHub issues. Additionally, our experiments on each step highlights the effectiveness of the strategies we proposed. Finally, a user study with 26 participants further confirms the high quality of our repairs. All our code and prompts are released at our GitHub repository  [^1] https://github.com/UGAIForge/DesignRepair.
@@
  • We systematically analyzed Material Design guidelines and converted them into structured knowledge bases, i.e., low-level component knowledge base and high-level system design knowledge base, enabling efficient and effective utilization of knowledge in design quality assurance processes.
  • We introduced DesignRepair, a novel dual-stream knowledge-driven, LLM-based method, uniquely designed to detect and repair design quality issues in frontend code. This approach simultaneously considers both the source code and the user-perceived rendered page, ensuring outstanding performance in design quality enhancement.
@@ II-A Material Design Guidelines
Material Design 3 [15] is a broad design system built and supported by Google designers and developers, and can be implemented across Android, iOS, and web platforms. Its latest version, Material Design 3, provides guidelines and examples to support important practices of user interface, ensuring personal, adaptive, and expressive experiences. It covers topics ranging from fine-grained design guidelines such as the application of individual components and color usage for specific meanings, to the high-level composition of components and common design patterns.
@@ III Approach
For the offline knowledge base construction phase (Fig. 3-A), we built a knowledge base (KB) consisting of two parts, i.e., a low-level Component Knowledge Base (KB-Comp) and a high-level System Design Knowledge Base (KB-System). This knowledge base functions as a domain expert, offering guidance for addressing potential UI issues. Given the frontend code and rendered page, we employ a parallel dual-pipe method to extract the used components and their corresponding property groups (as shown in Fig. 3-B). Finally, we implement a knowledge-driven, LLM-based repair method (Fig. 3-C). This approach allows us to carefully analyze and repair issues concurrently. By employing a divide and conquer strategy, we tackle each component/property group separately before integrating the repairs. This ensures a cohesive, optimized final output, achieved through a thorough and scrutinized repair process.
@@ III-A Knowledge Base Construction
As introduced in Section II-A, we utilize Material Design 3 as our foundational knowledge base. By systematically analyzing the guidelines for efficient knowledge retrieval, we restructured them into two parts: a low-level Component Knowledge Base (KB-Comp) and a high-level System Design Knowledge Base (KB-System). KB-Comp provides detailed guidance for atomic-level elements, while KB-System addresses the broader aspects of aesthetic and functional harmony in interface design. For example, a guideline from KB-Comp might ensure a button is sized appropriately for accessibility purpose, while a guideline from KB-System would focus on how this button fits into the overall page layout and flow. Within these knowledge bases, we further classified the information into soft constraints (i.e., recommended guidelines) and hard constraints (i.e., do/don’t guidelines), catering to different levels of design concerns. For instance, a soft constraint would suggest font styles that align with the brand’s aesthetic, whereas a hard constraint may specify the minimum text size for readability.
@@ III-A1 Component knowledge base
We created the component knowledge base (KB-Comp) at the component level, drawing from Components section in the Material Design (see Section II-A). Each component is linked with a range of guidelines. In total, we obtained 24 component types associating with Material Design guidelines. Fig. 4(a) illustrates the number of guidelines for each component type and the distribution of various UI/UX design aspects (like usage, responsive layout) that these guidelines address. The distribution of guidelines across these component types demonstrates a notable imbalance, showing a long-tail distribution pattern. For instance, buttons and navigation components are the most abundant, with 109 and 95 entities respectively, while checkboxes are the least represented, with only 6 entities. To facilitate understanding, Fig. 5 (a) shows two hard constraints and two soft constraints regarding the button component. In this example, the first hard constraint aligns with the guideline shown in Fig. 2(b), and the first soft constraint correlates with the guideline in Fig. 2(a).
@@
In conclusion, based on our detailed knowledge base, we propose knowledge-driven LLM-based method, which can effectively manage the long-tail distribution of design guidelines, ensuring an equitable application of these principles across all component types.
@@ III-A2 System Design knowledge base
To this end, we reformulated high-level design guidelines, i.e., Foundations and Styles sections (see Section II-A), into a structured system design knowledge base (KB-System). In total, We identified 12 system design aspects (e.g., typography), summarized 7 properties (e.g., clickable and spacing) and 15 mapping relations between them. For example, design aspects ‘role’ and ‘utilities’ in Styles - Color section, as well as the design aspect accessibility ‘contrast’ in Foundations - Color section are both connected to the ‘color’ property. These relations are crucial for marking guidelines, aiding LLMs in understanding the intricate relationships between design elements and their practical applications in a web development context. Based on this, we reconstructed the design guidelines into triples of $<$system design aspect, property, relation$>$. Fig. 5 (b) shows two examples of the system design knowledge base.
Foundations have 7 key aspects including Flow, Layout, Implementation, Structure, Label, Text, and Color. For instance, the ‘Flow’ aspect relates to how users navigate through an interface, while ‘Color’ pertains to both aesthetic appeal and accessibility. In total, we identified 118 knowledge entities for Foundations.
Styles consist of 5 key aspects: Typography, Elevation, Icons, Color, and Shape. These elements craft the UI’s visual identity, adaptable through Material themes. In our work, we structured 80 knowledge entities related to Styles, with Icons specifically linked to icon buttons and Shapes applicable across all component types.
Property Groups include 7 categories: Group, Clickable, Spacing, Platform, Label, Text, and Color. These are actionable property levels that we map to high-level guidelines for effective detection and repair.
@@ III-B1 Components Extraction
The first aspect of our strategy is to identify a list of component types within the Page Under Test (PUT) using GPT-4. This initial step is important for the efficient retrieval of relevant component knowledge from KB-Comp in subsequent repair stages.
@@ III-C1 Relevant Knowledge Retrieval
(a) Component Mapping and Knowledge Retrieval: To map extracted web page components to the component names in our KB-Comp for alignment, we use LLMs for reasoning with the prompt $P_{\text{map\_kb}}$. This approach allows us to handle the diversity in terminology used across different front-end libraries such as Shadcn[13], Tailwind CSS[14] without the need of defining any exhaustive rule-based mapping function. For instance, a ‘navigation bar’ component might be referred to as ‘Navbars’ in Tailwind CSS, while in Shadcn UI it could be labeled as ‘Navigation Menu’. Such variability in naming conventions can be effectively addressed by LLMs without the need to define an exhaustive mapping function. The mapping process is represented by:
@@ III-C3 Overall Repair
In this final step, we integrate the repair suggestions obtained from both the component-level and property-group analyses. This integration ensures a holistic repair of the frontend code.
Formally, given the original PUT code $In_{\text{PUT}}$ and the set of partially repaired code snippets $C3$, our objective is to use these inputs to produce the final, fully repaired page. To achieve this, we employ prompt with $P_{\text{all}}$, which is designed to direct the overall repair process:
@@
The $P_{\text{all}}$ prompt effectively merges the various repair suggestions, taking into account any potential overlaps or conflicts between the proposed fixes. The goal is to ensure that the final output, represented as Repaired_page, is not only free from identified design flaws but also aligns with the established design guidelines and usability standards. Due to the page limit, all prompts can be obtained at our GitHub repository.
@@ IV Experiments
  • RQ1: How effective is our approach in detecting violations according to Design Guidelines compared to the baseline?
  • RQ2: How effective are individual strategies and the overall repair approach?
  • RQ3: What is the perceived quality of the fixes generated by our approach from user perspectives?
@@ IV-A Subjects
We conducted two experiments to evaluate the performance of our frontend code repair approach on both generated code from Vercel’s V0 and real projects from GitHub.
@@ IV-A1 Vercel’s V0 Projects
Our research focuses on examining code and the rendered pages generated by LLM tools, specifically targeting adherence to design guidelines and enabling code-level adjustments. Our work aims to assist developers with limited knowledge of UI design guidelines who utilize LLM-generated UI mock-ups as starting points. Vercel’s V0 [9] generated projects were chosen for this purpose, as it is a leading LLM tool that attracted 100K users within three weeks of its release [31]. Additionally, Vercel’s V0 is widely used to assist in frontend development, with 259K results on GitHub as of December 2024.
@@ IV-A2 Github Projects
To further evaluate the effectiveness of our method in real-world projects, we searched for “top frontend projects” and “frontend example projects” on GitHub and Google, identifying the top 6 popular projects. These projects represent various real-world website uses, ranging from commercial product frontend websites to blogs, with star counts ranging from 1.1k to 5.6k. Their popularity and representativeness have proven their usefulness and quality.
We randomly selected 66 files out of 577 files from these projects, ensuring diverse functionalities. These projects utilize a variety of representative packages, including React [12], Next.js [32], Material-UI [33], and TypeScript [34], which can examine the capabilities of our DesignRepair in detecting and repairing design issues across these different technologies. Following the same protocol in Section IV-A1, we identified 1,793 components/property groups across these projects, of which 115 guidelines were violated.
@@ IV-A3 Subject Analysis
For Vercel’s V0 projects, there were 33.45 components/property group per file and 9.8 guideline violations per project. Of these, 149 (76%) were component-related and 47 (24%) were property-related. Additionally, 27 (14%) were soft constraint violations, while 169 (86%) were hard constraint violations. For GitHub projects, there were 27 components/property group per file and 1.74 guideline violations per project. Among these, 60 (52%) were component-related and 55 (48%) were property-related. Soft constraint violations accounted for 15 (13%), and hard constraint violations accounted for 100 (87%).
Based on our analysis, the generated frontend code from Vercel’s V0 was found to have complexity comparable to real GitHub projects (33.45 vs. 27 components/property group per file). This suggests the potential of LLM-generated code to match the structural richness required for real frontend development. However, GitHub projects have significantly fewer violations per project compared to Vercel’s V0 projects (1.74 vs. 9.8), highlighting a substantial gap in quality. This indicates that while LLMs can generate complex code, their outputs are prone to a higher rate of errors and require enhanced automated quality assurance mechanisms to improve code quality. In addition, Vercel’s V0 projects have a higher percentage of component-related issues (over 52%) compared to property-level issues. Real-world projects, although generally better designed, still have minor issues such as inconsistent padding or missing alt-text.
In conclusion, we found that human-generated issues differ significantly from AI-generated ones, making it valuable to analyze them separately for better insights.
IV-B RQ1: Effectiveness in detecting violations